An amateur astronomer decides to build a telescope from a discarded pair of eyeglasses. One of the lenses has a refractive power of 10 diopters, while the other has a refractive power of 1.5 diopters. (a) Which lens should be the objective? (b) How far apart should the lenses be separated? (c) What is the angular magnification of the telescope?
Answer #1

focal length of a lens =1/refractive power

the focal length of first lens f1 = 1/10=0.1m=10cm

the focal length of second lens f2 = 1/1.5=0.666m=66.6cm

larger focal length of lens is objective lens and the smaller is eye piece

a) so objective lens fo = 66.6 cm i.e., 1.6diopter lens

and eye piece fe = 10 cm

b) the distance between the lens L=fo+fe=66.6+10=76.6 cm

c) angular magnification m=fo/fe

m=76.6 / 10

m = 7.66

The angular Magnification is 7.66.

I hope help you !!